# Decomposed Quadratization: Efficient QUBO Formulation for Learning Bayesian Network

### v7

@@ 1 Introduction
In this study, we propose a QUBO formulation specifically designed to use fewer binary variables than quadratization. We demonstrate that this formulation offers significant advantages for learning Bayesian networks in terms of bit capacity. Figure 1 illustrates the overview of our approach. Experimental results on $16$ instances, ranging from $37$ to $223$ variables, show that our formulation dramatically reduces the number of binary variables compared to (O’Gorman et al. 2014). These instances can be encoded on the circuit of the Fujitsu Digital Annealer, a fully-coupled annealing machine with $100$K bit capacity (Nakayama et al. 2021). The scores achieved by the Digital Annealer using our formulation were greater than those of existing approaches.
@@ 3 Decomposed Quadratization
Let =$\mathcal{F}=\{\{1,2,3,4\},\{1,2,3\},\{4\},\{3,4\}\}$. A decomposed quadratization requires $3$ auxiliary variables: $Q_{1}=\{1,2,3\},Q_{2}=\{4\}$, and $Q_{3}=\{3\}$. A quadratization requires $6$ auxiliary variables: $Q_{1}=\{1\},Q_{2}=\{2\},Q_{3}=\{3\},Q_{4}=\{4\},Q_{5}=\{1,2\}$, and $Q_{6}=\{3,4\}$.
@@
1:  $W_{\mathrm{open}}\leftarrow\mathcal{F},V_{\mathrm{close}}\leftarrow\emptyset,V_{\mathrm{open}}\leftarrow\bigcup_{V\in\mathcal{F}}2^{V}\setminus\{\emptyset\}$.
@@
Figure 2: Example of the QUBO formulation with candidate parent subsets. An undirected edge between two binary variables represents that the objective function contains their product term. Gray signifies that the state of a binary variable is 00 and white is $1$. Each color corresponds to the edge in the DAG. Let $n=3,m=2,\lambda_{1}=3,\lambda_{2}=1,\lambda_{3}=1,\Pi_{1}=\{X_{2}\},\Pi_{2}=\emptyset,\Pi_{3}=\{X_{1},X_{2}\},W_{11}=\{X_{2}\},W_{12}=\{X_{3}\},W_{13}=\{X_{2},X_{3}\},W_{21}=\{X_{1},X_{3}\},W_{31}=\{X_{1},X_{2}\},\mu_{1}=2,\mu_{2}=1,\mu_{3}=1,U_{11}=\{X_{2}\},U_{12}=\{X_{3}\},U_{21}=\{X_{1},X_{3}\}$, and $U_{31}=\{X_{1},X_{2}\}$.
@@ 4 Efficient QUBO for Structure Learning
1:  $W_{\mathrm{open}}\leftarrow\mathcal{F}_{i},V_{\mathrm{close}}\leftarrow\emptyset,V_{\mathrm{open}}\leftarrow\bigcup_{V\in\mathcal{F}_{i}}2^{V}\setminus\{\emptyset\}$.
@@
4:     Execute the row from $2$ to $4$ in algorithm 1.
@@
[] Instance $m$ O’Gorman et al. IP only IP + algorithm 2 Number Time [s] Number Time [s] Number Time [s] alarm $4$ $3982\pm 346$ $272\pm 61$ $\bm{1373\pm 46}$ $61\pm 29$ $\bm{1373\pm 46}$ $\bm{14\pm 2}$ barley $4$ $1826\pm 279$ $271\pm 149$ $\bm{995\pm 143}$ $\bm{11\pm 2}$ $\bm{995\pm 143}$ $\bm{11\pm 2}$ hailfinder $4$ $2374\pm 31$ $52\pm 24$ $\bm{1913\pm 6}$ $\bm{4\pm 1}$ $\bm{1913\pm 6}$ $\bm{4\pm 1}$ hepar2 $4$ $3629\pm 423$ $51\pm 16$ $\bm{2616\pm 133}$ $3\pm 0$ $\bm{2616\pm 133}$ $\bm{2\pm 0}$ chess $2$ $7040\pm 167$ $\bm{63\pm 2}$ $\bm{6964\pm 165}$ $89\pm 4$ $\bm{6964\pm 165}$ $103\pm 4$ chess $3$ $80116\pm 3014$ $\bm{3229\pm 151}$ $26355\pm 1403$ $7268\pm 492$ $\bm{23662\pm 1103}$ $5567\pm 376$ win95pts $2$ $5150\pm 39$ $\bm{27\pm 2}$ $\bm{5028\pm 42}$ $33\pm 3$ $\bm{5028\pm 42}$ $37\pm 4$ win95pts $3$ $37705\pm 1935$ $1223\pm 90$ $9220\pm 398$ $1036\pm 117$ $\bm{9025\pm 333}$ $\bm{890\pm 105}$ pathfinder $2$ $10831\pm 382$ $\bm{178\pm 8}$ $\bm{10725\pm 379}$ $259\pm 30$ $\bm{10725\pm 379}$ $284\pm 20$ pathfinder $3$ $107725\pm 5329$ $\bm{11695\pm 834}$ $38245\pm 3242$ $30707\pm 4322$ $\bm{34448\pm 3016}$ $21364\pm 3773$ mushroom $2$ $18050\pm 508$ $\bm{318\pm 13}$ $\bm{17930\pm 506}$ $681\pm 32$ $\bm{17930\pm 506}$ $513\pm 23$ connect $2$ $15155\pm 334$ $\bm{137\pm 9}$ $\bm{14978\pm 340}$ $228\pm 17$ $\bm{14978\pm 340}$ $240\pm 12$ connect $3$ $179788\pm 6242$ $\bm{9897\pm 602}$ $64533\pm 4352$ $30381\pm 4267$ $\bm{58247\pm 4053}$ $20106\pm 2661$ munin1 $2$ $25180\pm 266$ $\bm{699\pm 20}$ $\bm{25001\pm 268}$ $953\pm 75$ $\bm{25001\pm 268}$ $1079\pm 67$ andes $2$ $29502\pm 171$ $\bm{40\pm 1}$ $\bm{28810\pm 182}$ $249\pm 37$ $\bm{28810\pm 182}$ $250\pm 35$ andes $3$ $92838\pm 3735$ $2182\pm 289$ $\bm{31162\pm 136}$ $504\pm 132$ $\bm{31162\pm 136}$ $\bm{217\pm 8}$
Table 1: Number of binary variables and execution time for the QUBO formulation. The mean and standard deviation of $5$ trials are presented. All the candidate parent subsets not proven optimal within $60$ [s] were feasible solutions. The bold values represent the smallest number of binary variables and the shortest execution times among all approaches.
@@ 5.1 QUBO Formulation
Evaluation.  The comparison between IP + algorithm 2 and IP only in table 1 demonstrates that optimizing candidate parent subsets with algorithm 2 is advantageous in terms of both the number of binary variables and execution time when the parent set size is larger. The quadratization for some instances with $m=3$ required binary variables over the $100$K bit capacity of the Digital Annealer. The decomposed quadratization in our approach significantly reduced the required number of binary variables compared to the quadratization approach. The execution time of our formulation was disadvantageous for most larger networks, primarily due to the increasing number of terms. Table 2 shows that incorporating the candidate parent subsets into the QUBO introduced a significant bottleneck.
@@ 5.2 Score Maximization
[] Instance $m$ Time [s] alarm $4$ $1\pm 0,3\pm 1,10\pm 1$ barley $4$ $0\pm 0,0\pm 0,11\pm 2$ hailfinder $4$ $0\pm 0,0\pm 0,4\pm 1$ hepar2 $4$ $0\pm 0,0\pm 0,2\pm 0$ chess $2$ $2\pm 0,0\pm 0,100\pm 4$ chess $3$ $406\pm 49,2554\pm 199,2608\pm 195$ win95pts $2$ $1\pm 0,0\pm 0,36\pm 4$ win95pts $3$ $45\pm 5,398\pm 82,446\pm 54$ pathfinder $2$ $3\pm 0,0\pm 0,281\pm 19$ pathfinder $3$ $1407\pm 272,1679\pm 184,18278\pm 3349$ mushroom $2$ $23\pm 1,0\pm 0,490\pm 22$ connect $2$ $5\pm 0,0\pm 0,235\pm 12$ connect $3$ $2472\pm 436,4129\pm 148,13505\pm 2114$ munin1 $2$ $12\pm 1,0\pm 0,1067\pm 67$ andes $2$ $1\pm 0,0\pm 0,249\pm 35$ andes $3$ $6\pm 1,2\pm 1,209\pm 8$
@@
[] Instance $m$ DAQ GOQ HCS SAO GOB alarm $4$ 00 00 00 00 $\bm{5}$ barley $4$ 00 00 00 00 $\bm{5}$ hailfinder $4$ 00 00 00 00 $\bm{5}$ hepar2 $4$ 00 00 00 00 $\bm{5}$ chess $2$ $\bm{5}$ 00 00 00 00 chess $3$ $\bm{5}$ 00 00 00 00 win95pts $2$ $1$ 00 00 00 $\bm{4}$ win95pts $3$ $2$ 00 00 $\bm{3}$ 00 pathfinder $2$ $\bm{4}$ 00 00 $1$ 00 pathfinder $3$ $\bm{5}$ 00 00 00 00 mushroom $2$ $\bm{5}$ 00 00 00 00 connect $2$ $\bm{5}$ 00 00 00 00 connect $3$ $\bm{2}$ $1$ $1$ $1$ 00 munin1 $2$ $\bm{5}$ 00 00 00 00 andes $2$ $\bm{5}$ 00 00 00 00 andes $3$ $\bm{5}$ 00 00 00 00
@@ D.2 Identification of Candidate Parent Sets
[] Instance $m$ Number Time [s] alarm $4$ $2291\pm 210$ $1700\pm 11$ barley $4$ $355\pm 12$ $235015\pm 95243$ hailfinder $4$ $692\pm 16$ $90247\pm 3334$ hepar2 $4$ $689\pm 59$ $40216\pm 534$ chess $2$ $32570\pm 1686$ $59\pm 0$ chess $3$ $236285\pm 15636$ $5463\pm 284$ win95pts $2$ $9231\pm 665$ $60\pm 3$ win95pts $3$ $43657\pm 3850$ $2404\pm 51$ pathfinder $2$ $50962\pm 3254$ $320\pm 5$ pathfinder $3$ $398131\pm 47889$ $44991\pm 3426$ mushroom $2$ $273526\pm 12647$ $689\pm 26$ connect $2$ $76996\pm 4381$ $331\pm 7$ connect $3$ $710684\pm 73964$ $83361\pm 6624$ munin1 $2$ $135570\pm 6765$ $2466\pm 29$ andes $2$ $7424\pm 219$ $852\pm 52$ andes $3$ $14195\pm 556$ $104683\pm 3855$
Table 4: Number of candidate parent sets and execution time for their identification. The mean and standard deviation of $5$ trials are presented. Empty sets were excluded from the candidate parent sets.
@@ D.3 Classical Simulated Annealing
3:     if a random number between 00 and $1$ is less than $\exp(-\Delta/T_{0}^{1-\mathrm{current~{}time}/\mathrm{time~{}limit}})$then
@@ D.5 Identification of Candidate Parent Subsets
[] Instance $m$ Time [s] alarm $4$ $1\pm 0,50\pm 28,10\pm 1$ barley $4$ $0\pm 0,1\pm 0,10\pm 2$ hailfinder $4$ $0\pm 0,0\pm 0,4\pm 1$ hepar2 $4$ $0\pm 0,0\pm 0,2\pm 0$ chess $2$ $6\pm 0,0\pm 0,83\pm 3$ chess $3$ $639\pm 58,3150\pm 167,3480\pm 315$ win95pts $2$ $2\pm 2,0\pm 0,31\pm 1$ win95pts $3$ $49\pm 8,530\pm 73,457\pm 66$ pathfinder $2$ $18\pm 2,0\pm 0,241\pm 29$ pathfinder $3$ $3844\pm 750,2132\pm 129,24731\pm 3458$ mushroom $2$ $277\pm 19,1\pm 0,403\pm 14$ connect $2$ $26\pm 3,0\pm 0,203\pm 15$ connect $3$ $6219\pm 1396,4568\pm 177,19595\pm 2743$ munin1 $2$ $78\pm 6,0\pm 0,875\pm 71$ andes $2$ $1\pm 0,0\pm 0,247\pm 37$ andes $3$ $4\pm 0,7\pm 1,493\pm 131$

### v8

@@ 1 Introduction
In this study, we propose a QUBO formulation specifically designed to use fewer binary variables than quadratization. We demonstrate that this formulation offers significant advantages for learning Bayesian networks in terms of bit capacity. Figure 1 illustrates the overview of our approach. Experimental results on $16$ instances, ranging from $37$ to $223$ variables, show that our formulation dramatically reduces the number of binary variables compared to (O’Gorman et al. 2014). These instances can be encoded on the circuit of the Fujitsu Digital Annealer, a fully-coupled annealing machine with 100100K bit capacity (Nakayama et al. 2021). The scores achieved by the Digital Annealer using our formulation were greater than those of existing approaches.
@@ 3 Decomposed Quadratization
Let $\mathcal{F}=\{\{1,2,3,4\},\{1,2,3\},\{4\},\{3,4\}\}$. A decomposed quadratization requires $3$ auxiliary variables: $Q_{1}=\{1,2,3\},Q_{2}=\{4\}$, and $Q_{3}=\{3\}$. A quadratization requires $6$ auxiliary variables: $Q_{1}=\{1\},Q_{2}=\{2\},Q_{3}=\{3\},Q_{4}=\{4\},Q_{5}=\{1,2\}$, and $Q_{6}=\{3,4\}$.
@@
[…]
@@
Figure 2: Example of the QUBO formulation with candidate parent subsets. An undirected edge between two binary variables represents that the objective function contains their product term. Gray signifies that the state of a binary variable is 0 and white is $1$. Each color corresponds to the edge in the DAG. Let $n=3,m=2,\lambda_{1}=3,\lambda_{2}=1,\lambda_{3}=1,\Pi_{1}=\{X_{2}\},\Pi_{2}=\emptyset,\Pi_{3}=\{X_{1},X_{2}\},W_{11}=\{X_{2}\},W_{12}=\{X_{3}\},W_{13}=\{X_{2},X_{3}\},W_{21}=\{X_{1},X_{3}\},W_{31}=\{X_{1},X_{2}\},\mu_{1}=2,\mu_{2}=1,\mu_{3}=1,U_{11}=\{X_{2}\},U_{12}=\{X_{3}\},U_{21}=\{X_{1},X_{3}\}$, and $U_{31}=\{X_{1},X_{2}\}$.
@@ 4 Efficient QUBO for Structure Learning
[…]
@@
4:  Execute the row from $2$ to 44 in algorithm 1.
@@
[…]
Table 1: Number of binary variables and execution time for the QUBO formulation. The mean and standard deviation of 55 trials are presented. All the candidate parent subsets not proven optimal within $60$ [s] were feasible solutions. The bold values represent the smallest number of binary variables and the shortest execution times among all approaches.
@@ 5.1 QUBO Formulation
Evaluation.  The comparison between IP + algorithm 2 and IP only in table 1 demonstrates that optimizing candidate parent subsets with algorithm 2 is advantageous in terms of both the number of binary variables and execution time when the parent set size is larger. The quadratization for some instances with $m=3$ required binary variables over the 100100K bit capacity of the Digital Annealer. The decomposed quadratization in our approach significantly reduced the required number of binary variables compared to the quadratization approach. The execution time of our formulation was disadvantageous for most larger networks, primarily due to the increasing number of terms. Table 2 shows that incorporating the candidate parent subsets into the QUBO introduced a significant bottleneck.
@@ 5.2 Score Maximization
[] Instance $m$ Time [s] alarm 44 $1\pm 0,3\pm 1,10\pm 1$ barley 44 $0\pm 0,0\pm 0,11\pm 2$ hailfinder 44 $0\pm 0,0\pm 0,4\pm 1$ hepar2 44 $0\pm 0,0\pm 0,2\pm 0$ chess 22 $2\pm 0,0\pm 0,100\pm 4$ chess 33 $406\pm 49,2554\pm 199,2608\pm 195$ win95pts 22 $1\pm 0,0\pm 0,36\pm 4$ win95pts 33 $45\pm 5,398\pm 82,446\pm 54$ pathfinder 22 $3\pm 0,0\pm 0,281\pm 19$ pathfinder 33 $1407\pm 272,1679\pm 184,18278\pm 3349$ mushroom 22 $23\pm 1,0\pm 0,490\pm 22$ connect 22 $5\pm 0,0\pm 0,235\pm 12$ connect 33 $2472\pm 436,4129\pm 148,13505\pm 2114$ munin1 22 $12\pm 1,0\pm 0,1067\pm 67$ andes 22 $1\pm 0,0\pm 0,249\pm 35$ andes 33 $6\pm 1,2\pm 1,209\pm 8$
@@
[] Instance $m$ DAQ GOQ HCS SAO GOB alarm 44 0 0 0 0 $\bm{5}$ barley 44 0 0 0 0 $\bm{5}$ hailfinder 44 0 0 0 0 $\bm{5}$ hepar2 44 0 0 0 0 $\bm{5}$ chess 22 $\bm{5}$ 0 0 0 0 chess 33 $\bm{5}$ 0 0 0 0 win95pts 22 11 0 0 0 $\bm{4}$ win95pts 33 22 0 0 $\bm{3}$ 0 pathfinder 22 $\bm{4}$ 0 0 11 0 pathfinder 33 $\bm{5}$ 0 0 0 0 mushroom 22 $\bm{5}$ 0 0 0 0 connect 22 $\bm{5}$ 0 0 0 0 connect 33 $\bm{2}$ 11 11 11 0 munin1 22 $\bm{5}$ 0 0 0 0 andes 22 $\bm{5}$ 0 0 0 0 andes 33 $\bm{5}$ 0 0 0 0
@@ D.2 Identification of Candidate Parent Sets
[] Instance $m$ Number Time [s] alarm 44 $2291\pm 210$ $1700\pm 11$ barley 44 $355\pm 12$ $235015\pm 95243$ hailfinder 44 $692\pm 16$ $90247\pm 3334$ hepar2 44 $689\pm 59$ $40216\pm 534$ chess 22 $32570\pm 1686$ $59\pm 0$ chess 33 $236285\pm 15636$ $5463\pm 284$ win95pts 22 $9231\pm 665$ $60\pm 3$ win95pts 33 $43657\pm 3850$ $2404\pm 51$ pathfinder 22 $50962\pm 3254$ $320\pm 5$ pathfinder 33 $398131\pm 47889$ $44991\pm 3426$ mushroom 22 $273526\pm 12647$ $689\pm 26$ connect 22 $76996\pm 4381$ $331\pm 7$ connect 33 $710684\pm 73964$ $83361\pm 6624$ munin1 22 $135570\pm 6765$ $2466\pm 29$ andes 22 $7424\pm 219$ $852\pm 52$ andes 33 $14195\pm 556$ $104683\pm 3855$
Table 4: Number of candidate parent sets and execution time for their identification. The mean and standard deviation of 55 trials are presented. Empty sets were excluded from the candidate parent sets.
@@ D.3 Classical Simulated Annealing
3:  if a random number between 0 and $1$ is less than $\exp(-\Delta/T_{0}^{1-\mathrm{current~time}/\mathrm{time~limit}})$then
@@ D.5 Identification of Candidate Parent Subsets
[] Instance $m$ Time [s] alarm 44 $1\pm 0,50\pm 28,10\pm 1$ barley 44 $0\pm 0,1\pm 0,10\pm 2$ hailfinder 44 $0\pm 0,0\pm 0,4\pm 1$ hepar2 44 $0\pm 0,0\pm 0,2\pm 0$ chess 22 $6\pm 0,0\pm 0,83\pm 3$ chess 33 $639\pm 58,3150\pm 167,3480\pm 315$ win95pts 22 $2\pm 2,0\pm 0,31\pm 1$ win95pts 33 $49\pm 8,530\pm 73,457\pm 66$ pathfinder 22 $18\pm 2,0\pm 0,241\pm 29$ pathfinder 33 $3844\pm 750,2132\pm 129,24731\pm 3458$ mushroom 22 $277\pm 19,1\pm 0,403\pm 14$ connect 22 $26\pm 3,0\pm 0,203\pm 15$ connect 33 $6219\pm 1396,4568\pm 177,19595\pm 2743$ munin1 22 $78\pm 6,0\pm 0,875\pm 71$ andes 22 $1\pm 0,0\pm 0,247\pm 37$ andes 33 $4\pm 0,7\pm 1,493\pm 131$